# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ Abstract
The rapid advancements in artificial intelligence (AI), particularly in generative AI and large language models (LLMs), have profoundly impacted the creative industries by enabling innovative content creation, enhancing workflows, and democratizing access to creative tools. This paper explores the significant technological shifts since our previous review in 2022, highlighting how these developments have expanded creative opportunities and efficiency. These technological advancements have enhanced the capabilities of text-to-image, text-to-video, and multimodal generation technologies. In particular, key breakthroughs in LLMs have established new benchmarks in conversational AI, while advancements in image generators have revolutionized content creation. We also discuss AI integration into post-production workflows, which has significantly accelerated and refined traditional processes. Despite these innovations, challenges remain—particularly for the media industry—due to the demands on communication traffic from creative content. We therefore include data compression and quality assessment in this paper. Furthermore, we highlight the trend toward unified AI frameworks capable of addressing multiple creative tasks and underscore the importance of human oversight to mitigate AI-generated inaccuracies. Finally, we explore AI’s future potential in the creative sector, stressing the need to navigate emerging challenges to maximize its benefits while addressing associated risks.
@@ 1 Introduction
Artificial intelligence (AI) has grown dramatically over the past few years, particularly due to the rise of generative AI and large language models (LLMs). These advancements have been widely regarded as beneficial in many countries, such as the UK, as highlighted in a report published by the Authority of the House of Lords [1]. These AI advancements have also had significant direct and indirect impacts on the creative industries, influencing the direction of their growth. Generative AI, for instance, primarily focuses on generating new data that is not identical to the training data but shares similarities. However, the training data can be very large and have a much broader range than what humans have encountered themselves. The resulting output may present entirely new work to an individual, serving as inspiration. Indeed, AI tools have also opened up opportunities to a wider range of users with different skill sets, enabling creative work to be done faster, more effectively, and with greater creativity. Moreover, these new technologies not only influence artists and creative industries but also open new ways for audiences to experience arts and culture [2].
@@
The major breakthrough in generative AI has come from OpenAI [^1] https://openai.com/, an AI research and deployment company, with the introduction of Generative Pre-trained Transformer (GPT) models for LLMs. LLMs are specifically designed to understand and generate human language. They are characterized by their vast size in terms of parameters and the amount of training data. This breakthrough was particularly significant with the release of ChatGPT in 2022, which was fine-tuned from a model in the GPT-3.5 series. ChatGPT is a conversational model that includes advancements in safety features to mitigate the generation of inappropriate content. Several other LLM platforms were developed around the same time, such as LaMDA and PaLM by Google AI, Ernie Bot by Baidu, and BLOOM by BigScience. Additionally, Anthropic launched Claude, the LLM trained specifically to be harmless and honest, leveraging reinforcement learning from human feedback (RLHF) - a technique used to train AI systems to appear more human [3]. However, ChatGPT stands out as the most renowned, thanks to its quick and efficient responses, and notably, it is available for free. Another breakthrough in the same year was in text-to-image models. OpenAI achieved a significant milestone with DALL·E 2, producing impressive artworks and photorealistic images despite its limited language understanding. Midjourney by Midjourney, Inc., another well-known text-to-image generation, supports higher resolution images, which can go up to 4096$\times$4096 pixels. Stable Diffusion by Stability AI, for which the code and model weights are publicly available [^2] https://github.com/Stability-AI/stablediffusion, allows developers and artists to further adapt it to their own applications.
@@
The next breakthrough happened in 2023 when OpenAI unveiled GPT-4, a significantly larger model with more parameters and improved performance compared to its predecessors [4]. GPT-4 is a multimodal large language model that can generate responses to both text and images. It incorporates DALL·E 3, enabling it to comprehend a much broader range of nuances and details than earlier versions. In March 2024, Claude 3 Opus by Anthropic has been released, boasting multimodal capabilities in generating images, tables, graphs and diagrams. Moreover, Anthropic claims that Claude 3 Opus outperforms GPT-4 in generating human-like dialog and contextually aware responses. These AI advancements however make the media industry faces challenges. With the impact of these technologies, DMG Media, the Financial Times, and the Guardian Media Group have highlighted concerns about the potential impact on print journalism, particularly if AI tools reduce the need for users to click through to news websites, affecting advertising and subscription revenues [1].
@@
Generating videos is significantly more challenging than generating images. In February 2024, Google announced Gemini 1.5 with a capability of processing a large data (approximately 8 times more than GPT-4), like video footage and audio recordings [^3] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/. In the same month, OpenAI provided the first preview of Sora, a model capable of generating impressive realistic videos up to 1 minute long. As of November 2024, Sora remains unreleased and unavailable to the public. Based on the videos released by OpenAI, Sora appears to outperform other text-to-video models. A month later, Gemini 1.5 has announced its support native audio understanding in 180$+$ countries. With the emergence of Gemini 1.5 and Sora, alongside other forthcoming AI tools, it is obvious that video content creation will significantly benefit, opening up the media landscape for creativity and providing more opportunities for diverse storytellers, while also saving time to produce. A recent example is the AI-generated Christmas commercial by Coca-Cola [^4] https://www.youtube.com/watch?v=4RSTupbfGog. These advertisements overcome the limitations of current technologies by using very short videos with quick scene transitions, ensuring that any artifacts, such as unnatural fingers, are not immediately noticeable.
@@
For post-production workflow, generative AI may not have a direct impact, but the neural networks originally proposed for generative AI have been widely adapted to serve this purpose, leading to significant improvements in both output quality and computational speed. Moreover, in recent years, there has been a noticeable trend in technology towards adopting a unified framework rather than addressing individual challenges, as it better reflects real-world scenarios. For instance, natural history filmmaking involves challenging acquisition environments and high production standards. It is often filmed in low light, underwater, heat haze, or adverse weather conditions. This can lead to increased noise levels, focus issues, low contrast, color balance problems, and blurriness in the footage. Moreover, the unified models offer advantages in generalizing to diverse tasks and providing flexibility. Take Painter by BAAI Vision [5] as an example, which employs an image pair as a task prompt (similar to a text prompt in LLMs), their model transfers the input image to produce a similar output as the task prompt, enabling it to undertake various tasks such as segmentation, low-light enhancement, rain removal, etc.
@@
While generative AI makes creating and post-processing digital content easier and faster, more efficient performance in video content streaming is consequently required. Although AI-based solutions have been proposed both for conventional coding tool enhancement and for new compression frameworks, they have not been fully adopted in practical applications due to hardware constraints and complexity issues. However, the latest learning-based video codecs have already demonstrated their potential to compete with conventional standard video codecs, in particular, with the latest advances in implicit neural representations. The techniques could emerge in the next few years and contribute to the future video coding standards.
@@
Moreover, in the past few years, we have also seen the impact of AI in the research area of visual quality assessment. The advances include new model architectures based on different attention mechanisms and the application of LLMs, which evidently improves model generalization. New training methodologies have also been proposed based on weakly/unsupervised learning, which address the issue with limited training content.
@@
One of the exciting aspects of using LLMs in the creative sector is that ‘The human in the loop’ [6] is simplified through text prompts, as language enables artists to convey complex emotions and narratives. This is important because generative AI could produce mistakes, known as hallucinations; hence, we, as humans, correct this through reinforcement learning with human feedback [7].
@@
In this paper, the objective is to investigate the latest advancements in technology that have emerged since our previous review paper on AI in the creative industries before 2022 [8]. Compared to the earlier paper, where most technologies were developed as support tools, this updated review paper reveals significant shifts in the technologies after 2022 due to generative AI and most recent AI-based technologies. We explore the new direction of how AI could be involved in the creative sector. Similar to [8], we first provide a high-level overview of current advanced AI technologies (Section 2), followed by a selection of creative domain applications (Section 3), where current technologies enabling AI in creative scenarios are described. Finally, we discuss challenges and the future potential of AI associated with the creative industries (Section 4).
@@ 2 Current Advanced AI Technologies
Foundation models (FMs) have been termed by The Stanford Institute for Human-Centered Artificial Intelligence in 2021 [9] which mean “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. LLMs is one kind of the FMs.
As this paper provides updates from our previous review [8], an introduction to AI, basic neurons, convolutional neural networks (CNNs), generative adversarial networks (GANs), recurrent neural networks (RNNs), and deep reinforcement learning (DRL) can be found there. Here, we will emphasize four technologies that are rapidly growing and largely affecting creative industries, particularly during the year around 2023. This includes Transformers, Large language models (LLMs), Diffusion Models (DMs), and Implicit Neural Representations (INRs). It is important to note that while these newer technologies are gaining prominence, those from previous generations are still widely used, often in conjunction with the newer ones, as they offer distinct advantages. For instance, CNNs effectively capture local features and semantic meaning, while the attention mechanism in transformers offers global dependencies.
@@ 2.1 Transformers
No objection, the usage of AI has become widely accessible to everyone due to the launch of ChatGPT by OpenAI since November, 2022. ChatGPT has become the fastest-growing consumer software application in history [^5] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/, marking a significant advancement in the performance of large language models (LLMs). However, the turning point was actually back in 2017 when Google AI introduced ‘Transformer’ architectures in their publication ‘Attention Is All You Need’ [10]; since then, the performance of language models has dramatically improved. The applications have been extended, including vision understanding [11], and even learning multiple tasks across multiple modalities simultaneously (e.g., Gato [12]). Here, we will provide a background of transformers.
Before the advent of transformers, natural language processing (NLP) was done using recurrent neural networks (RNNs), processing data sequences sequentially. In contrast, the transformers capture long-range dependencies through self-attention mechanisms, that go across all words in the sequence, weigh the importance of different words and find relationships between words regardless of their positions. This learns context-aware representation and enables parallel processing of the entire sequence, making the transformers computationally efficient. A set of several attention layers running in parallel is called Multi-Head Attention. In the technical overview level, Transformer architecture, shown in Fig. 1 (a), comprises Encoder and Decoder, similar to many CNN-based generators. But the encoder is a stack of identical layers, concatenating a multi-head self-attention mechanism and a fully connected feed-forward network. The decoder is also a stack of identical layers, of which each layer has additional sub-layer to perform multi-head attention over the output of the encoder stack.
@@
Attention modules have not only been using in the transformers, but they have also been integrated with other deep learning architectures, including with CNNs in particular, used for image classification [13], object detection [14], and other computer vision tasks [15] as providing performance improvement.
In 2020, the first successful training of a transformer encoder for image recognition was published in [11], known as Vision Transformer (ViT). ViT divides an input image into patches, similar to words in a sentence, and processes them through multi-head attention. Additionally, a Multilayer Perceptron (MLP) is employed as the feedforward network. The later work uses a hierarchical division of image inputs and the shifted window approach, called Swin Transformer [16] by Microsoft. It reported outperforming ViT by 2.4% in ImageNet-22K classification (21,841 different categories). Its version 2 [17] applies cosine function in the attention module enabling the scaling up of capacity and resolution. More detail on transformer-based object detection is discussed in Section 3.4.2. To date, Swin Transformer has been widely adopted in various applications, including image restoration [18], where it is integrated with UNet. The application of transformers to video recognition has done by input as a sequence of frames [19] or dividing the video into space-time blocks [20]. Comprehensive surveys on transformers for images and videos can be found in [21] and [22], respectively.
Transformers have been widely used and offer better performance across many tasks. One reason for this is the availability of the open-source Transformers library through Hugging Face [^6] https://huggingface.co/, a platform that assists developers in building their own applications across various tasks, including computer vision, NLP, audio, tabular data, multimodal tasks, and reinforcement learning. The platform also provides model zoo (pretrained networks) and datasets.
Note that in recent years, emerging technologies have adopted state space models [23, 24], commonly known as ‘Mamba.’ These models are recognized as a linear variant of Transformers and are distinguished by their linear complexity in attention modeling. It is widely reported that they offer performance on par with, or even surpassing, traditional Transformers, while demanding fewer computational resources and less memory.
Figure 1: Generative AI. (a) Transformer architecture [10]. (b) The top row is diffusion process and the bottom row is generation process of new image [25]. (c) Latent Diffusion Models (LDM) [26].

### v3

@@ Abstract
The rapid advancements in artificial intelligence (AI), particularly in generative AI and large language models (LLMs), have profoundly impacted the creative industries, enabling more innovative content creation, enhancing workflows, and democratizing access to creative tools. This paper explores these technological shifts, with particular focus on how those that have emerged since our previous review in 2022 have expanded creative opportunities and improved efficiency. These technological advancements have enhanced the capabilities of text-to-image, text-to-video, and multimodal generation technologies. In particular, key breakthroughs in LLMs have established new benchmarks in conversational AI, while advancements in image generators have revolutionized content creation. We also discuss the integration of AI into post-production workflows, which has significantly accelerated and improved traditional processes. Once content has been created, it must be delivered to its audiences the media industry is facing the demands of increased communication traffic due to creative content. We therefore include a discussion of how AI is beginning to transform the way we represent and compress media content. We highlight the trend toward unified AI frameworks capable of addressing and integrating multiple creative tasks, and we underscore the importance of human insight to drive the creative process and oversight to mitigate AI-generated inaccuracies. Finally, we explore AI’s future potential in the creative sector, stressing the need to navigate emerging challenges and to maximize its benefits while addressing the associated risks.
@@ 1 Introduction
The influence of artificial intelligence (AI) has grown dramatically over the past few years, particularly due to the rise of generative AI and large language models (LLMs). These advancements are widely regarded as beneficial by many countries, creating significant opportunities for growth (e.g. as outlined in the UK, by the Authority of the House of Lords [1]). These advances have also had significant direct and indirect impacts on the creative industries, influencing the direction of their growth. Generative AI, for instance, primarily focuses on generating new data that is not identical to the training data yet shares similarities with it. However, the cardinality of the training data can be huge, larger than what any individual human has ever encountered. The resulting output may therefore act as a new source of insipration.
AI tools also provide opportunities for a wider range of users to work more efficiently and effectively, with even greater creativity. Moreover, these new technologies not only influence creators, but they also enable new ways for audiences to experience art and culture [2].
@@
A major breakthrough in generative AI has been led by OpenAI [^1] https://openai.com/, an AI research and deployment company, with their introduction of Generative Pre-trained Transformer (GPT) models for LLMs. LLMs are specifically designed to understand and generate human language. They are characterized by their vast size in terms of parameters and the amount of training data used to create them. This breakthrough was particularly impactful when the company released ChatGPT in 2022, which was fine-tuned from a model in the GPT-3.5 series. ChatGPT is a conversational model that includes advanced safety features that mitigate the generation of inappropriate content. Several other LLM platforms were also developed contemporaneously, such as LaMDA and PaLM by Google AI, Ernie Bot by Baidu, and BLOOM by BigScience. Additionally, Anthropic launched Claude, the LLM trained specifically to be harmless and honest, leveraging reinforcement learning from human feedback (RLHF) - a technique used to train AI systems to appear more human [3]. Nonetheless, ChatGPT stands out as the most renowned, thanks to its quick and efficient responses, and notably its public accessibility, being available for free.
Another breakthrough in 2022 was in the area of text-to-image models. OpenAI achieved a significant milestone with DALL·E 2, producing impressive artworks and photorealistic images despite its limited language understanding. Midjourney by Midjourney, Inc., another well-known text-to-image generator, supports higher resolution images, up to 4096$\times$4096 pixels. Stable Diffusion by Stability AI, for which the code and model weights are publicly available [^2] https://github.com/Stability-AI/stablediffusion, allows developers and artists to further adapt AI to suit their own specific applications.
@@
The next breakthrough happened in 2023 when OpenAI unveiled GPT-4, a significantly larger model with estimated 1.8 trillion parameters and improved performance compared to its predecessors [4]. However, this still represents less than 1% of the human brain’s approximately 600 trillion synaptic connections [^3] https://www.rsb.org.uk/biologist-features/ai-versus-the-brain. GPT-4 is a multimodal large language model that can generate responses to both text and images. It incorporates DALL·E 3, enabling it to comprehend a much broader range of nuances and details than earlier versions. In March 2024, Claude 3 Opus by Anthropic was released, boasting multimodal capabilities in generating images, tables, graphs and diagrams. Moreover, Anthropic claims that Claude 3 Opus outperforms GPT-4 in generating human-like dialog and contextually aware responses. These rapid advances have, in turn, led the creative industries to face significant challenges. For example, DMG Media, the Financial Times, and Guardian Media Group have highlighted concerns about the potential impact on print journalism, particularly if AI tools reduce the need for users to click through to news websites, affecting advertising and subscription revenues [1]. There is also concern about ‘AI-generated slop’—low-quality, mass-produced content created by AI that often lacks coherence or originality [^4] https://reutersinstitute.politics.ox.ac.uk/news/ai-generated-slop-quietly-conquering-internet-it-threat-journalism-or-problem-will-fix-itself. It is typically used for spam, search engines, or clickbait, and is criticized for cluttering the internet and undermining genuine human-created content.
@@
The generation of videos is significantly more challenging for AI than generating images. In February 2024, Google announced Gemini 1.5 which had the capability to process approximately 8 times more data than GPT-4, opening opportunities for video and audio processing  [^5] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/. In the same month, OpenAI provided its first preview of Sora, a model capable of generating impressive realistic videos up to 1 minute long. Based on the videos released by OpenAI, Sora appears to outperform other text-to-video models. Sora is currently available to ChatGPT subscribers. A month later, Gemini 1.5 announced its support for native audio understanding in 180$+$ countries. With the emergence of these tools, together with the prospect of further advances, it is clear that video content creation will be a major beneficiary. This will further open up the media landscape for creativity and provide more opportunities for diverse storytellers, while also reducing production time. A recent example is the AI-generated Christmas commercial by Coca-Cola [^6] https://www.youtube.com/watch?v=4RSTupbfGog. Such advertisements overcome the limitations of current technologies by using very short videos with rapid scene transitions, ensuring that any artifacts, such as unnatural fingers, are less apparent.
@@
For the case of post-production workflows, generative AI may not have a direct impact, but the neural networks originally proposed for generative AI have been widely adapted to serve this purpose. This has led to significant improvements in both output quality and computational speed. Moreover, there is a noticeable trend towards adopting a unified framework rather than addressing individual tasks, as it better reflects real-world scenarios. For instance, natural history filmmaking involves challenging acquisition environments and high production standards. Filming often takes place in low light conditions, in the presence of heat haze, underwater or in adverse weather conditions. This often results in increased noise levels, focus issues, low contrast, color balance problems, and blurriness in the footage. In such cases, Unified models can offer advantages in generalizing to diverse tasks and providing flexibility. Take Painter by BAAI Vision [5] as an example, which employs an image pair as a task prompt (similar to a text prompt in LLMs), their model transfers the input image to produce a similar output as the task prompt, enabling it to undertake various tasks such as segmentation, low-light enhancement or rain removal.
@@
While generative AI can facilitate and accelerate the creation and post-processing of digital media, there is an equivalent need to transmit or stream it efficiently to users. Although AI-based solutions have been proposed both for the enhancement of conventional video coding tools and for new compression frameworks, they are yet to be adopted in practical applications due to hardware constraints, complexity issues and a lack of standardization. Despite this, the latest learning-based video codecs have already demonstrated their potential to compete with conventional standardized video codecs and are being actively investigated in various standards bodies such as MPEG and AOM.
@@
Furthermore, in recent years, AI has also impacted our ability to assess and monitor the perceptual quality of visual media. Advances have included new model architectures based on different attention mechanisms and the application of LLMs, which evidently improve model generalization. New training methodologies have also been proposed based on weakly/unsupervised learning, which address issues associated with the limited availability of labeled training content.
@@
One of the exciting aspects of using LLMs in the creative sector is that ‘The human in the loop’ [6] is simplified through text prompts, with sophisticated, multilingual language capabilities enabling artists to convey complex emotions and narratives. This is important because generative AI does produce mistakes, known as hallucinations. Human oversight is thus essential to correct this through reinforcement learning with feedback [7].
@@
In this paper, the objective is to reveal to the reader, the latest technology advancements that have emerged since our previous review paper on AI in the creative industries (published in 2022) [8]. Compared to this earlier paper, which was written when most AI technologies were used as support tools, this updated review describes thesignificant disruptive shifts that have emerged over the past 3-4 years driven by generative AI and other recent AI-based technologies. Similar to [8], we first provide a high-level overview of current advanced AI technologies (Section 2), followed by a selection of key creative domain applications (Section 3) where current AI technologies are changing creative practice. Finally, we discuss the challenges and the future potential of AI associated with the creative industries (Section 4).
@@ 2 Current Advanced AI Technologies
This paper provides a review of AI in the creative industries, building on our previous publication in 2022 [8]. The reader is referred to that work for an introduction to AI, basic neurons, convolutional neural networks (CNNs), generative adversarial networks (GANs), recurrent neural networks (RNNs) and deep reinforcement learning (DRL). In this paper, we emphasize four key technologies that have grown in importance since 2022 that have had a significant impact on the creative industries. These are Transformers, Large language models (LLMs), Diffusion Models (DMs), and Implicit Neural Representations (INRs). It is important to note that, while these newer technologies are gaining prominence, those from previous generations remain in widespread use, often in conjunction with the newer ones. For instance, CNNs complement transformers since CNNs effectively capture local features and semantic meaning, while the attention mechanism in transformers capture global dependencies.
One important class within AI that has become dominant since our previous review comprises Foundation models (FMs). These were described by The Stanford Institute for Human-Centered Artificial Intelligence in 2021 [9] as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. Foundation models have been enabled by rapid advances in AI-oriented computing power and have been underpinned the emergence and success of Large Language Models, particularly following the launch of ChatGPT by OpenAI in 2022. ChatGPT has become the fastest-growing consumer software application in history [^7] https://www.reuters.com/technology/chatgpt-sets-record-fastest-growing-user-base-analyst-note-2023-02-01/.
These technologies are expanded on below.
@@ 2.1 Transformers
In 2017, Google AI introduced the concept of ‘Transformer’ architectures in their publication ‘Attention Is All You Need’ [10]. This work has since, been instrumental in the development and success of large language models alongside many other applications, including vision understanding [11], and multiple modality learning (e.g., Gato [12]).
Before the advent of transformers, natural language processing (NLP) was performed using recurrent neural networks (RNNs), processing data sequences sequentially. In contrast, the ability of transformers to capture long-range dependencies through self-attention mechanisms that extend across all words in the sequence, meant that the importance of different words could be established globally, understanding relationships regardless of their positions. This context-aware representation enables parallel processing of the entire sequence, making the transformers computationally efficient. A set of several attention layers running in parallel is called Multi-Head Attention.
The Transformer architecture, shown in Fig. 1 (a), comprises Encoder and Decoder sections, similar to many CNN-based generators. However the encoder is now a stack of identical layers, concatenating a multi-head self-attention mechanism and a fully connected feed-forward network. The decoder is also a stack of identical layers, in which each layer has additional sub-layer to perform multi-head attention over the output of the encoder stack.
@@
It should be noted that attention modules are not solely used in transformers, but have also been successfully integrated into other deep learning architectures such as CNNs, used for image classification [13], object detection [14], and other computer vision tasks [15].
In 2020, the first successful training of a transformer encoder for image recognition was published [11], reeferred to as a Vision Transformer (ViT). The ViT decomposes an input image into patches, similar to words in a sentence, and processes them through multi-head attention. Additionally, a Multilayer Perceptron (MLP) is employed as the feedforward network. In later work Microsoft introduced a hierarchical division of image inputs and a shifted window approach in their Swin Transformer [16]. This was reported to outperform ViT by 2.4% in ImageNet-22K classification (21,841 different categories). Its version 2 [17] applied a cosine function in the attention module. enabling the scaling of capacity and resolution. More detail on transformer-based object detection is discussed in Section 3.4.2. To date, Swin Transformers have been widely adopted in a range of applications including image restoration [18].
Comprehensive surveys on the use of transformers for image and video processing can be found in [19] and [20], respectively.
Transformers have been widely used and offer better performance across many tasks. One reason for this widespread adoption has been the availability of open-source Transformer libraries such as Hugging Face [^8] https://huggingface.co/, a platform that assists developers to build applications for tasks including computer vision, NLP, audio, tabular data, multimodal tasks, and reinforcement learning. The platform also provides access to model zoo  [^9] Such as https://modelzoo.co/ and https://pytorch.org/serve/model_zoo.html pretrained networks and datasets.
In recent years state space models [21, 22], commonly known as ‘Mamba’ have emerged. These are a linear variant of Transformersdistinguished by their linear complexity in attention modeling. They are acknowledged to offer an equivalent or better performance than traditional Transformers, while demanding fewer computational resources and less memory.
Figure 1: Generative AI. (a) Transformer architecture [10]. (b) The top row represents the diffusion process and the bottom row represents the generation process of the new image [23]. (c) Latent Diffusion Models (LDM) [24].